Grade 12th passAlgebra

Express each of the following as a product of prime factors only in exponential form:
729 x 64

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

The factors of 729 = 3 × 3 × 3 × 3 × 3 × 3
=36
The factors of 64 = 2 × 2 × 2 × 2 × 2 × 2
= 26
Then,
= (36 × 26)
= 36 × 26

​Thanks